Find all n ? N, such that the equation x-8/n-10=n/x has no solution x.

Grade:10

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
(x-8)/(n-10)=n/x
=>x^2-8x-n(n-10)=0
No solution
=>64+4(1)n(n-10)<0
=>16+n^2-10n<0
=>(n-8)(n-2)<0
=>2<n<8

=>n=3,4,5,6,7
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty